OpenStudy (lukecrayonz):

(8/z)-z=4 (8/z)-z=4 @Mathematics

OpenStudy (a_clan):

multiply by z throughout

OpenStudy (lukecrayonz):

Not sure what you mean...

OpenStudy (lukecrayonz):

I got it guys :)

OpenStudy (anonymous):

(8/z)-z=4 => 8-z^2=4z =>z^2+4z=8 =>z(z+4)=8 either z=8 or z=4

OpenStudy (lukecrayonz):

Oh, haha :PP